[figure: 9.7 cm]

CuSiO_{3}: A Candidate System for Purely Oxygen Antiferromagnet?

M. Baenitz^{a}, S.-L. Drechsler^{b}, A.A. Gippius^{c}, E.N. Morozova^{c}, A.S. Moskvin^{d}

aMax Planck Institut für Chemische Physik Fester Stoffe, D-01187, Dresden, Germany bInstitut für Festkörper- und Werkstofforschung Dresden, D-01171, Dresden, Germany

cMoscow State University, 119899, Moscow, Russia dUral State University, 620083, Ekaterinburg, Russia

Contrary to CuGeO_{3} with its spin-Peierls transition, the isostructural CuSiO_{3} system seems to be typical S = 1/2 antiferromagnet with long range AFM ordering below T_{N} ≈ 8.0 K. NQR measurements in CuSiO_{3} have revealed an unexpected puzzling picture evidencing the absence of the noticeable local hyperfine magnetic field on ^{63,65}Cu nuclei in AFM ordered state, and considerable charge redistribution as compared with CuGeO_{3} counterpart system. The ^{63,65}Cu-NQR spectrum in CuSiO_{3} system is likely to coincide with that in diamagnetic Cu_{2}O with the non-magnetic Cu^{1+} ionic state of copper. This coincidence may be accidental, however, small linewidth and weak temperature dependence both for NQR frequency and linewidth between 4.2 and 50 K without any anomaly at T_{N} are surprising. Moreover, the low-temperature behavior of T_{1}^{−1} in CuSiO_{3} reveals the non-magnetic quadrupole mechanism of spin-lattice relaxation. Such a behavior of copper ions in CuSiO_{3} points to a nonmagnetic Cu^{1+} copper valent state with the antiferromagnetically ordered oxygen surrounding. It means that the hole in the CuO_{2} cell should be localized at oxygen, rather than at copper sites. We suggest a qualitative physical reasoning for the ”sp hybridization induced oxygen hole stabilization effect” which is believed to be more pronounced in CuSiO_{3} as compared with CuGeO_{3}.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : A.S. Moskvin

Address for correspondence : Lenin str. 51, Ural State University, 620083, Ekaterinburg, Russia

Email address : alexandr.moskvin@usu.ru